OpenStudy (anonymous):

what is the radical of 60?

OpenStudy (amistre64):

what?

OpenStudy (amistre64):

\[\sqrt{60}\]by chance?

OpenStudy (amistre64):

4*15 so pull out the sqrt(4)

OpenStudy (amistre64):

if we arereducing or simplifying that is

OpenStudy (anonymous):

yes to simplest terms is it 4 square 15

OpenStudy (anonymous):

i want it broken down to its simplest form

OpenStudy (amistre64):

sqrt4 * sqrt15

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

48?

OpenStudy (amistre64):

48 what?

OpenStudy (anonymous):

what is the radical of 48

OpenStudy (amistre64):

sqrt(48) = sqrt(4*12) = sqrt(4*4*3) = sqrt(16*3) = sqrt(16) * sqrt(3)

OpenStudy (anonymous):

so its 16 sqrt 3

OpenStudy (anonymous):

or 3 sqrt 16